OpenStudy (anonymous):

y = (x - 3 )^2 +4 please help solve by identify the minimum or maximum y value

OpenStudy (whpalmer4):

That's a parabola. If you multiply it out, you'll see that the coefficient on the \(x^2\) term is positive, which means the parabola opens up (looks like a bowl). That means the vertex is the minimum point. You conveniently have the formula in what is called vertex form: \[y =a (x-h)^2+k\] where the vertex is at \((h,k)\). \(a = 1\) in this case. Where is the vertex?

OpenStudy (whpalmer4):

If you had the formula in expanded or standard form: \[y = ax^2+bx+c\]you can find the x coordinate of the vertex by \[x = -\frac{b}{2a}\]and plug that value of \(x\) into the formula to find the y coordinate of the vertex. Don't need to know that to do this problem, but it beats factoring to get vertex form if you don't already have it.

OpenStudy (anonymous):

I am not sure

OpenStudy (lncognlto):

The x^2 value is always going to be positive, as the parabola opens upwards and thus will have a minimum value, so it would need to equal 0 for the minimum y value.

OpenStudy (whpalmer4):

Compare your equation with mine: \[y = (x-3)^2+4\]\[y=(x-h)^2+k\] what are the values of \(h\) and \(k\)?

OpenStudy (anonymous):

please give me a moment to write out

OpenStudy (whpalmer4):

sure, I'm not going anywhere...

OpenStudy (anonymous):

I don't know I am having a hard time figuring out algebra

OpenStudy (anonymous):

I have a tutor that sits with me once a week an I just can seem to grab on to algebra

OpenStudy (anonymous):

are you there?

OpenStudy (whpalmer4):

yes, just got back.

OpenStudy (whpalmer4):

\[y=(x-h)^2+k\]\[y=(x-3)^2+4\] Look at the corresponding spots in each equation. Let's find the value of \(k\) first. There are 3 "chunks" in the second equation: \[y\]\[(x-3)^2\]\[4\] which of those most resembles \(k\) in the first equation? \(4\), right?

OpenStudy (anonymous):

yes

OpenStudy (whpalmer4):

Okay, now let's find \(h\). Any ideas?

OpenStudy (anonymous):

h would be 3 right?

OpenStudy (whpalmer4):

yes! (sorry about the delay, my family is bugging me for help on other projects)

OpenStudy (anonymous):

understand

OpenStudy (whpalmer4):

So our vertex is at \((h,k)\) and as we found that \(h = 3, k = 4\), that means the vertex is at \((3,4)\)

OpenStudy (whpalmer4):

There's a nice page talking about parabolas and their formulas here: http://www.mathwarehouse.com/geometry/parabola/standard-and-vertex-form.php

OpenStudy (anonymous):

okay thank you